# CMHSU: An R Statistical Software Package to Detect Mental Health Status, Substance Use Status, and their Concurrent Status in the North American Healthcare Administrative Databases

### v2

@@ Abstract
The concept of concurrent mental health and substance use (MHSU) and its detection in patients has garnered growing interest among psychiatrists and healthcare policymakers over the past four decades. Researchers have proposed various diagnostic methods, including the Data-Driven Diagnostic Method (DDDM), for the identification of MHSU. However, the absence of a standalone statistical software package to facilitate DDDM for large healthcare administrative databases has remained a significant gap. This paper introduces the R statistical software package CMHSU, available on the Comprehensive R Archive Network (CRAN), for the diagnosis of mental health (MH), substance use (SU), and their concurrent status (MHSU). The package implements DDDM using hospital and medical service physician visit counts along with maximum time span parameters for MH, SU, and MHSU diagnoses. A working example with a simulated real-world dataset is presented to explore three critical dimensions of MHSU detection based on the DDDM. Additionally, the limitations of the CMHSU package and potential directions for its future extension are discussed.
@@ 1.1 Concurrent Mental Health Status and Substance Use Status: Definition & Diagnosis
The concept of concurrent mental health (MH) and substance use (SU) disorders, collectively referred to as MHSU, has been explored for over four decades [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], with significant attention given to its methodological challenges [14]. Various approaches have been developed to diagnose MHSU, each with distinct applications and strengths. The primary diagnostic methods are outlined below:
@@ 1.1.5 Data-Driven Diagnostic Method (DDDM)
The DDDM leverages clinical and administrative data analysis to identify co-occurring mental health and substance use disorders through the utilization of healthcare databases and ICD codes. This approach is based on patterns of healthcare usage, such as hospitalizations and physician visits [23, 24, 25, 26]. For example, a healthcare researcher may analyze hospital records to identify individuals who, within the past year, had at least one hospitalization or two medical physician visits for mental health and substance use disorders, as indicated by ICD-10 codes [24, 25, 26].
@@ 1.2 North American Healthcare Administrative Databases
Country Level Hospital Data Physician Data Canada Federal Discharge Abstract Database (DAD) National Ambulatory Care Reporting System (NACRS) CIHI Physician Claims Data Canadian MIS Database (CMDB) Province MED-ÉCHO (Quebec) data Ontario Health data Alberta Health data Medical Services Plan (MSP) Data (BC) Alberta Health Physician Claims Régie de l’assurance maladie du Québec (RAMQ) Ontario Health Insurance Plan (OHIP) Territory Yukon Hospital Corporation Dis-charge Data Northwest Territories Health Authority Records Yukon Health Care Insurance Plan Claims Northwest Territories Medical Billing Data United States Federal National Inpatient Sample (NIS) Medicare Provider Analysis and Review (MEDPAR) Veterans Affairs (VA) Inpatient Data Medicare Physician/Supplier Part B Claims Data National Ambulatory Medical Care Survey (NAMCS) Veterans Affairs (VA) Outpatient Data State State Inpatient Databases (SID) California Office of Statewide Health Planning and Development (OSHPD) New York Statewide Planning and Research Cooperative System (SPARCS) State Medicaid Claims Data All-Payer Claims Databases (APCDs: e.g., Massachusetts, Colorado) Territory Guam Memorial Hospital Authority Records American Samoa Dept. of Health Inpatient Data Guam Medicaid and CHIP Claims Data American Samoa Medicaid Claims1
@@ 1.3 Motivation
    Scalability and Large-Scale Analysis: The Data-Driven Diagnostic Method (DDDM) offers the advantage of scalability by utilizing healthcare databases and administrative records to analyze large populations. Unlike approaches that require direct patient interaction, DDDM enables the comprehensive assessment of co-occurring mental health and substance use disorders across diverse demographic groups and healthcare systems [23].
@@
    Population-Level Insights for Policy and Planning: Through its capacity to analyze large-scale healthcare data, DDDM facilitates the identification of utilization trends, disparities, and service gaps. This allows policymakers to design targeted interventions and optimize resource allocation to address the needs of individuals with co-occurring disorders [24].
@@ 1.4 Study Outline
This paper is structured as follows: Initially, we provide essential theories and instructions for the installation of the CMHSU package. We introduce the four primary functions of the package, which are designed to detect mental health (MH), substance use (SU), and their concurrent status (MHSU) in both basic and comprehensive forms. Subsequently, we describe a simulation study using a simulated real-world healthcare administrative dataset to identify MH, SU, and MHSU statuses. This study includes three sub-studies where certain parameters from Definition 1.1 are held constant while others vary, allowing us to examine trends in the detection of MH, SU, and MHSU statuses. We conclude with a discussion on the contributions of the CMHSU package to existing literature, its limitations, and future development plans.
@@ 2.1 The Background and Installation
The CMHSU package utilizes the Data-Driven Diagnostic Method (DDDM) detailed in Section 1.1 and illustrated in Figure 1. It requires four R packages: dplyr [34], tidyr [35], purrr [36], and magrittr [37], and was developed using R version 4.2.2 [38]. The package (version 0.0.6.9) is accessible on the Comprehensive R Archive Network (CRAN) at
@@ 2.2 Detection of Mental Health Status
As an initial working example, we utilize a simulated real-world dataset, as described in Section 3.1, to identify all patients with a mental health status corresponding to one of the following disorders: psychotic, mood, anxiety, or neurocognitive disorders. The detection criteria include at least one hospital visit or at least one medical service physician visit, within a maximum time span of two months (60 days). The corresponding R code for implementing this detection is shown below; it generates the results depicted in Figure 3a:
@@ 2.3 Detection of Substance Use Status
As the second working example, we identify all patients with a substance use status corresponding to one of the following disorders: alcohol use, fentanyl use, cannabis use, or cocaine use. The detection criteria require at least one hospital visit or one medical service physician visit, within a maximum allowable time span of two months (60 days). Below is the corresponding R code; it generates the results illustrated in Figure 3b:
@@ 2.4 Detection of Concurrent Mental Health & Substance Use (MHSU) Status – Part (I)
As the third working example, we build on the earlier two to detect the Concurrent Mental Health and Substance Use (MHSU) status for clients diagnosed with psychotic, mood, anxiety, or neurocognitive disorders in combination with alcohol, fentanyl, cannabis, or cocaine consumption-related disorders. We require a minimum of one hospital visit and one physician visit within two months (60 days) for each of MH and SU; the maximum time span between MH and SU statuses is set to one year (365 days). It is important to note that the total dataset time span is 363 days, which is slightly shorter than the specified maximum span of 365 days. The R code is below; results appear in Figure 3c:
@@ 3.1 Simulated Real-World Data
Group Size Visit Date Span Visit Length #Hospital #Physician SU (freq) MH (freq) Other (freq) 1 10 01.01.2024–31.01.2024 1 month 1 2 F100 (1) F060 (2) NA 2 20 01.02.2024–31.03.2024 2 months 2 4 T4041 (2) F063 (4) J10 (4) 3 30 01.04.2024–31.06.2024 3 months 3 6 F120 (3) F064 (6) I10 (3) 4 40 01.07.2024–31.12.2024 6 months 6 12 F140 (6) F067 (12) I10 (6), J10 (12) 5 25 01.11.2024–31.12.2024 2 months 3 6 F100 (3) NA J10 (6) 6 25 01.11.2024–31.12.2024 2 months 2 4 NA F060 (4) I10 (2) 7 50 01.11.2024–31.12.2024 2 months 1 2 NA NA I10 (1), J10 (2)
Notes: F100(Alcohol); F060(Psychotic); T4041(Fentanyl); F063(Mood); F120(Cannabis); F064(Anxiety); F140(Cocaine); F067(Neurocognitive); I10(Hypertension); J10(Influenza); NA(Not Applicable).
@@ 4.1 Summary and Contributions
1. Dimension of Time Span within MH and SU Diagnoses (Section 3.2): The time span within MH and SU diagnoses plays a critical role. For in-stance, if a maximum time span of 7 days is considered, only 48 out of 100 patients (48.8%) are captured, whereas extending the time span to 56 days captures 90 out of 100 patients (90.0%). This delicate situation raises the question: ”What is the appropriate maximum time span for MH and SU diagnoses?”
2. Dimension of Required Number of Visits (Section 3.3): The number of required hospital and medical service physician visits significantly influences the detection rates. For example, requiring two hospital visits and four physician visits captures 90 out of 100 patients (90.0%), while increasing the requirement to three hospital visits and six physician visits reduces the capture rate to 70 out of 100 patients (70.0%). This delicate balance prompts the question: ”What is the optimal number (or ratio) of visits needed?”
3. Dimension of Time Span for Concurrent Status (Section 3.4): The maximum time span between MH and SU diagnoses also plays a vital role. For instance, setting this span to one month captures 61 out of 100 patients (61.0%), while extending it to three months captures 84 out of 100 patients (84.0%). This scenario raises the question: ”What is the ideal maximum time span between MH and SU diagnoses?”
@@
4.2 Limitations and Future Work
@@
This work has several limitations, each of which presents opportunities for future extensions. First, the package currently generates outputs as dataframes only, without providing summary frequency statistics for mental health, sub-stance use, and their concurrent statuses. Calculating these statistics requires additional R programming. Second, the window lag in the function MHSU_status_broad() is fixed at one day (as illustrated in Figure 2). Adding a parameter to specify the length of this window lag would enhance the flexibility of the detection process for researchers. Finally, for large-scale databases, it may be more efficient to partition the input dataset into multiple disjoint subsets and apply the MHSU_status_broad() function to each subset separately to manage the extensive outputs. Introducing an additional parameter within the function to automate the partitioning of the input data ($inputdata$) would further optimize the computational process.
@@ 5 Conclusions
Supplementary Material. The following supporting information can be downloaded at: https://cran.r-project.org/package=CMHSU, CMHSU R package materials.

### v3

@@ Abstract
The concept of concurrent mental health and substance use (MHSU) and its detection in patients has garnered growing interest among psychiatrists and healthcare policymakers over the past four decades. Researchers have proposed various diagnostic methods, including the Data-Driven Diagnostic Method (DDDM), for the identification of MHSU. However, the absence of a standalone statistical software package to facilitate DDDM for large healthcare administrative databases has remained a significant gap. This paper introduces the R statistical software package CMHSU, available on the Comprehensive R Archive Network (CRAN), for the diagnosis of mental health (MH), substance use (SU), and their concurrent status (MHSU). The package implements DDDM using hospital and medical service physician visit counts along with maximum time span parameters for MH, SU, and MHSU diagnoses. A working example using a simulated real-world dataset is presented to examine various analytical aspects, including three key dimensions of MHSU detection based on the DDDM framework, as well as temporal analysis to demonstrate the package’s application for healthcare policymakers. Additionally, the limitations of the CMHSU package and potential directions for its future extension are discussed.
@@ 1.1 Concurrent Mental Health Status and Substance Use Status: Definition & Diagnosis
The concept of concurrent mental health (MH) and substance use (SU) status, collectively referred to as MHSU, has been explored for over four decades [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], with significant attention given to its methodological challenges [14]. Various approaches have been developed to diagnose MHSU, each with distinct applications and strengths. The primary diagnostic methods are outlined below:
@@ 1.1.5 Data-Driven Diagnostic Method (DDDM)
The DDDM leverages clinical and administrative data analysis to identify co-occurring mental health and substance use status through the utilization of healthcare databases and the International Classification of Disease (ICD) codes. This approach is based on patterns of healthcare usage, such as hospitalizations and physician visits [23, 24, 25, 26]. For example, a healthcare researcher may analyze hospital records to identify individuals who, within the past year, had at least one hospitalization or two medical physician visits for mental health and substance use status, as indicated by ICD-10 codes [24, 25, 26].
@@ 1.2 North American Healthcare Administrative Databases
Country Level Hospital Data Physician Data Canada Federal Discharge Abstract Database (DAD) National Ambulatory Care Reporting System (NACRS) Canadian Institute of Health Information(CIHI) Physician Claims Data Canadian Management Information Database (CMDB) Province MED-ÉCHO (Quebec) data Ontario Health data Alberta Health data Medical Services Plan (MSP) Data (BC) Alberta Health Physician Claims Régie de l’assurance maladie du Québec (RAMQ) Ontario Health Insurance Plan (OHIP) Territory Yukon Hospital Corporation Dis-charge Data Northwest Territories Health Authority Records Yukon Health Care Insurance Plan Claims Northwest Territories Medical Billing Data United States Federal National Inpatient Sample (NIS) Medicare Provider Analysis and Review (MEDPAR) Veterans Affairs (VA) Inpatient Data Medicare Physician/Supplier Part B Claims Data National Ambulatory Medical Care Survey (NAMCS) Veterans Affairs (VA) Outpatient Data State State Inpatient Databases (SID) California Office of Statewide Health Planning and Development (OSHPD) New York Statewide Planning and Research Cooperative System (SPARCS) State Medicaid Claims Data All-Payer Claims Databases (APCDs: e.g., Massachusetts, Colorado) Territory Guam Memorial Hospital Authority Records American Samoa Dept. of Health Inpatient Data Guam Medicaid and CHIP Claims Data American Samoa Medicaid Claims1
@@
Remark 1.4.
The ICD codes provided in the following examples for mental health (MH) and substance use (SU) statuses are intended as illustrative examples rather than exhaustive lists. The complete set of ICD codes for these conditions may vary based on local and national guidelines, as well as the criteria established by study principal investigators.
@@ 1.3 Motivation
    Scalability and Large-Scale Analysis: The Data-Driven Diagnostic Method (DDDM) offers the advantage of scalability by utilizing healthcare databases and administrative records to analyze large populations. Unlike approaches that require direct patient interaction, DDDM enables the comprehensive assessment of co-occurring mental health and substance use status across diverse demographic groups and healthcare systems [23].
@@
    Population-Level Insights for Policy and Planning: Through its capacity to analyze large-scale healthcare data, DDDM facilitates the identification of utilization trends, disparities, and service gaps. This allows policymakers to design targeted interventions and optimize resource allocation to address the needs of individuals with co-occurring status [24].
@@ 1.4 Study Outline
This paper is structured as follows: Initially, we provide essential theories and instructions for the installation of the CMHSU package. We introduce the four primary functions of the package, which are designed to detect mental health (MH), substance use (SU), and their concurrent status (MHSU) in both basic and comprehensive forms. Subsequently, we describe a simulation study using a simulated real-world healthcare administrative dataset to identify MH, SU, and MHSU statuses. This study comprises four sub-studies. In the first three, specific parameters from Definition 1.1 are held constant while others vary, enabling an analysis of trends in the detection of MH, SU, and MHSU statuses. The final sub-study focuses on temporal analysis, demonstrating an example of package application for North American healthcare policymakers. We conclude with a discussion on the contributions of the CMHSU package to existing literature, its limitations, and future development plans.
@@ 2.1 The Background and Installation
The CMHSU package utilizes the Data-Driven Diagnostic Method (DDDM) parametric Definition 1.1 detailed in Section 1.1 and illustrated in Figure 1. It requires four R packages: dplyr [34], tidyr [35], purrr [36], and magrittr [37], and was developed using R version 4.2.2 [38]. The package (version 0.0.6.9) is accessible on the Comprehensive R Archive Network (CRAN) at
@@ 2.2 Detection of Mental Health Status
As an initial working example, we utilize a simulated real-world dataset, as described in Section 3.1, to identify all patients with a mental health status corresponding to one of the following status: psychotic, mood, anxiety, or neurocognitive disorders. The detection criteria include at least one hospital visit or at least one medical service physician visit, within a maximum time span of two months (60 days). The corresponding R code for implementing this detection is shown below; it generates the results depicted in Figure 3a:
@@ 2.3 Detection of Substance Use Status
As the second working example, we identify all patients with a substance use status corresponding to one of the following status: alcohol use, fentanyl use, cannabis use, or cocaine use. The detection criteria require at least one hospital visit or one medical service physician visit, within a maximum allowable time span of two months (60 days). Below is the corresponding R code; it generates the results illustrated in Figure 3b:
@@ 2.4 Detection of Concurrent Mental Health & Substance Use (MHSU) Status – Part (I)
As the third working example, we build on the earlier two to detect the Concurrent Mental Health and Substance Use (MHSU) status for clients diagnosed with psychotic, mood, anxiety, or neurocognitive disorders in combination with alcohol, fentanyl, cannabis, or cocaine consumption-related status. We require a minimum of one hospital visit and one physician visit within two months (60 days) for each of MH and SU; the maximum time span between MH and SU statuses is set to one year (365 days). It is important to note that the total dataset time span is 363 days, which is slightly shorter than the specified maximum span of 365 days. The R code is below; results appear in Figure 3c:
@@ 3.1 Simulated Real-World Data
Group Size Visit Date Span Visit Length #Hospital #Physician SU (freq) MH (freq) Other (freq) 1 10 01.01.2024–31.01.2024 1 month 1 2 F100 (1) F060 (2) NA 2 20 01.02.2024–31.03.2024 2 months 2 4 T4041 (2) F063 (4) J10 (4) 3 30 01.04.2024–31.06.2024 3 months 3 6 F120 (3) F064 (6) I10 (3) 4 40 01.07.2024–31.12.2024 6 months 6 12 F140 (6) F067 (12) I10 (6), J10 (12) 5 25 01.11.2024–31.12.2024 2 months 3 6 F100 (3) NA J10 (6) 6 25 01.11.2024–31.12.2024 2 months 2 4 NA F060 (4) I10 (2) 7 50 01.11.2024–31.12.2024 2 months 1 2 NA NA I10 (1), J10 (2) Notes: F100(Alcohol); F060(Psychotic); T4041(Fentanyl); F063(Mood); F120(Cannabis); F064(Anxiety); F140(Cocaine); F067(Neurocognitive); I10(Hypertension); J10(Influenza); NA(Not Applicable).
@@ 3.4 The Impact of Maximum Time Span for Concurrent Diagnosis
3.5 Temporal Analysis
In this section, we assume that, unlike the previous three sections (3.2-3.4), the study’s principal investigators have established agreed-upon default values for key parameters. These include a minimum of one hospital visit, a minimum of two medical service physician visits, a maximum time span of one month for mental health diagnoses, a maximum time span of one month for substance use diagnoses, and a maximum time span of one month for concurrent diagnoses. Here, policymakers are particularly interested in monitoring the monthly trends in the frequency and rate of MH, SU, and MHSU diagnoses throughout the 2024 calendar year. Figure 7 presents the results using similar programming presented in Appendix (Sections “A1. Scalability with large databases” and “A2. Summary Statistics Output”). As shown in the figure, both MH and SU exhibit an overall increasing trend over time, reaching their peak in November. In contrast, MHSU follows a fluctuating pattern initially, followed by a period of stabilization until November. However, all three categories experience a sharp decline in December.
Remark 3.1.
The temporal analysis presented here is based on frequency statistics. A similar approach applies to rate statistics, depending on the objectives of the principal investigators. The key difference lies in using proportions instead of counts in the estimation process, as outlined in Appendix (Section ”A2. Summary Statistics Output”).
@@ 3.5 Temporal Analysis
Figure 7: Monthly Frequency of MH, SU, and MHSU status in the calendar year 2024.
@@ 4.1 Summary and Contributions
  1. 1.
    Dimension of Time Span within MH and SU Diagnoses (Section 3.2): The time span within MH and SU diagnoses plays a critical role. For in-stance, if a maximum time span of 7 days is considered, only 48 out of 100 patients (48.8%) are captured, whereas extending the time span to 56 days captures 90 out of 100 patients (90.0%). This delicate situation raises the question: ”What is the appropriate maximum time span for MH and SU diagnoses?”
  2. 2.
    Dimension of Required Number of Visits (Section 3.3): The number of required hospital and medical service physician visits significantly influences the detection rates. For example, requiring two hospital visits and four physician visits captures 90 out of 100 patients (90.0%), while increasing the requirement to three hospital visits and six physician visits reduces the capture rate to 70 out of 100 patients (70.0%). This delicate balance prompts the question: ”What is the optimal number (or ratio) of visits needed?”
  3. 3.
    Dimension of Time Span for Concurrent Status (Section 3.4): The maximum time span between MH and SU diagnoses also plays a vital role. For instance, setting this span to one month captures 61 out of 100 patients (61.0%), while extending it to three months captures 84 out of 100 patients (84.0%). This scenario raises the question: ”What is the ideal maximum time span between MH and SU diagnoses?”
@@
4.2 Strengths, Limitations and Future Work
The R package CMHSU offers several unique features that make it a valuable and advantageous statistical tool for detecting mental health (MH), substance use (SU), and concurrent mental health and substance use (MHSU) status within large administrative healthcare databases. These features include:
  1. 1.
    Flexibility: CMHSU incorporates four core functions and ten customizable parameters, allowing researchers to account for a wide range of predefined scenarios by investigators when identifying MH, SU, and MHSU status in healthcare administrative databases.
  2. 2.
    Comprehensiveness: CMHSU enables detection of almost all mental health and substance use conditions than other statistical tools given their recent comprehensive codifications in ICD-10 and ICD-11.
  3. 3.
    Efficiency: The CMHSU package is designed for ease of use, requiring only a basic statistical background. Unlike advanced statistical and machine learning-based detection methods for MH[39], SU[40], and MHSU[41]— which typically necessitate expertise in topics such as K-Nearest Neighbors, Random Forests, Gradient Boosted Trees, Deep Neural Networks— CMHSU offers a user-friendly and time-efficient alternative, making it particularly accessible to researchers and practitioners without extensive statistical or machine learning expertise.
  4. 4.
    Interpretability: CMHSU employs a trace-back methodology to identify MH, SU, and MHSU status based on their corresponding ICD codes. This approach enhances the clarity and transparency of the results, facilitating faster interpretation, improved visualization of detected cases, and analysis of trends over time compared to other statistical tools.
  5. 5.
    Seamless Integration: The package is free and easy to install, ensuring compatibility with existing analytical tools used for processing large healthcare administrative databases. This seamless integration enhances its accessibility and usability in real-world research applications.
This work has several limitations, each of which presents opportunities for future extensions. Some of them are as follows:
  1. 1.
@@ 4.2 Strengths, Limitations and Future Work
    Scalability: For large-scale databases, it may be more efficient to partition the input dataset into multiple disjoint subsets and apply the MHSU_status_broad() function to each subset separately to manage the extensive outputs. Introducing an additional parameter within the function to automate the partitioning of the input data ($inputdata$) would further optimize the computational process(See Appendix “A1. Scalability with large databases” ).
  2. 2.
    Output Format: The package currently generates outputs as dataframes only, without providing summary frequency statistics for mental health, substance use, and their concurrent statuses. Calculating these statistics requires additional R programming (See Appendix “A2. Summary Statistics Output”).
  3. 3.
    Customization: The window lag in the function MHSU_status_broad() is fixed at one day (as illustrated in Figure 2). Adding a fixed or adaptive parameter to specify the length of this window lag would enhance the flexibility of the detection process for researchers. The choice of fixed or adaptive status depends to specific patient data in the study.
  4. 4.
    Evaluation: The DDDM approach assumes that ICD coding in administrative databases is reliable. However, in practice, these records are often incomplete or subject to misclassification, leading to potential biases in the evaluation process and, consequently, affecting the accuracy of the results. Furthermore, it is still unclear how to measure the package’s precision in detecting patients with MH status, SU status, and MHSU status, as well as how to compare its performance using DDDM with more advanced machine learning-based methods [39, 40, 41]. This comparison needs a mapping mechanism between the above parameters in the DDDM and above machine learning-based methods requiring subsequent methodological research.
  5. 5.
    Standardization: The use of ICD codes for diagnosis depends on local and national jurisdictions (Table 1). Potential inconsistencies in coding across different jurisdictions and healthcare databases may affect the reliability of the results for comparisons across these jurisdictions.
  6. 6.
    Empirical Validation: The simulation study in this paper uses self-generated simulated data. The optimal scenario would involve using a real-life empirical healthcare administrative database, which would significantly enhance the validity and usefulness of the results. However, key obstacles such as (1) Data Privacy and Security Regulations and (2) Access Restrictions and Bureaucratic Hurdles made it impossible to conduct the current analysis on such a database.
  7. 7.
    Geographical Adaptability: CMHSU is designed for detecting mental health status (MH), substance use status (SU), and their concurrent status (MHSU) within North American healthcare databases, based on the original DDDM developed by researchers in this region. However, it remains unclear whether these methods can be effectively adapted for use in non-North American healthcare systems, given potential differences in healthcare infrastructure, coding practices, and administrative data structures.
  8. 8.
    Temporal Analysis: The simulation study in this paper serves as a hypothetical demonstration of the package’s application. Its results can only be truly meaningful and applicable to real-world policymaking if based on a universally agreed-upon set of predefined parameters related to time span and visit definitions. However, no such universally accepted standard currently exists.
@@ 5 Conclusions
Supplementary Material.
S1. R Package Materials
The following supporting information can be downloaded at “CMHSU” R package CRAN documentation: https://cran.r-project.org/package=CMHSU.
S2. Shiny Application
A Shiny application has been included in the supplementary materials to facilitate interactive exploration of the estimations. The application can be accessed at: https://mohsensoltanifar.shinyapps.io/rpackageshinyappcmhsu/.
@@
Appendix
A1. Scalability with large databases :
The scalability of the large healthcare administrative dataset, inputdata, is achieved by considering two dimensions: unique IDs and time. First, for the unique ID dimension, the function splitfunction_id(), introduced below, takes the dataset inputdata (containing $m$ unique patients) and an integer $n\geq 1$ as input. It then partitions the dataset into $k$ subsets, where $k=[m/n]+1(if\ \ n\nmid m),m/n(if\ \ n\mid m).$ The first $k-1$ subsets each contain $n$ unique patients, while the kth subset contains at most $n$ unique patients. For example, we consider the case where $m=200,n=18$ and $k=12:$
line #1:> splitfunction_id <- function(inputdata, n) {                              inputdata <- inputdata %>arrange(ClientID)                              unique_ids <- unique(inputdata\$ClientID)                              groups <- split(unique_ids, ceiling(seq_along(unique_ids)/n))                              split_datasets <- list()                              for (i in seq_along(groups)) {                              split_datasets[[paste0("inputdata_", i)]] <- inputdata %>%                              filter(ClientID %ingroups[[i]])                              }                              return(split_datasets)                              }line #2:> inputdata_split_id <- splitfunction_id(SampleRWD, 18)line #3:> inputdata_split_id\$inputdata_1
Second, for the time dimension, the function splitfunction_time(), introduced below, takes the dataset inputdata (with time span $T$ days) and an integer $t\geq 1$ as input. It then partitions the dataset into $l$ subsets, where $l=[T/t]+1(if\ \ t\nmid T),T/t(if\ \ t\mid T).$ The first $l-1$ subsets each are in the time span of $t$ days, while the lth subset time span is at most $t$. For example, we consider the case where $T=363,t=30.5$ and $l=12:$
line #1:> splitfunction_time <- function(inputdata, t) {                                inputdata <- inputdata %>mutate(VisitDate = as.Date(VisitDate))                                inputdata <- inputdata %>arrange(VisitDate)                                VisitDate_0 <- min(inputdata\$VisitDate, na.rm = TRUE)                                split_datasets <- list()                                i <- 1                                while(TRUE) {                                start_date <- VisitDate_0 + (i - 1) * (t + 1)                                end_date <- start_date + t                                subset_data <- inputdata %>%                                filter(VisitDate >= start_date & VisitDate <= end_date)                                if (nrow(subset_data) == 0) break                                split_datasets[[paste0("inputdata_", i)]] <- subset_data                                i <- i + 1                                }                                return(split_datasets)                                }line #2:> inputdata_split_time <- splitfunction_time(SampleRWD, 30.5)line #3:> inputdata_split_time\$inputdata_1
A2. Summary Statistics Output:
The following R program computes summary statistics for MH, SU, and MHSU frequencies from the CMHSU output shown in the Figure 3 (c):
line #1:> myexample<-SampleRWD[,c(1:4)]line #2:> SampleMHSU_1 = MHSU_status_basic(myexample, n_MHH=1, n_MHP=1, n_SUH=1,                 n_SUP=1, t_MH=60, t_SU=60, t_MHSU=365,                 ICD_MH=c("F060","F063","F064","F067"),                 ICD_SU=c("F100","T4041","F120","F140"))line #3:> SummarySampleMHSU_1 <-  SampleMHSU_1 %>%          summarise(            MH_Count = sum(MH_status == "YES"),            MH_Proportion = formatC(mean(MH_status == "YES"), format = "f", digits = 3),            SU_Count = sum(SU_status == "YES"),            SU_Proportion = formatC(mean(SU_status == "YES"), format = "f", digits = 3),            MHSU_Count = sum(MHSU_status == "YES"),            MHSU_Proportion = formatC(mean(MHSU_status == "YES"), format = "f", digits = 3) )line #4:> print(SummarySampleMHSU_1)line #5:> # A tibble: 1 × 6          MH_Count MH_Proportion SU_Count SU_Proportion MHSU_Count MHSU_Proportion          <int>     <chr>        <int>     <chr>       <int>        <chr>          125       0.625         125      0.625        100         0.500